rg819616:

If 6 newborn babies are randomly selected, how many different gender sequences are possible? a.720 c.12 b.36 d.64

Pixel:

6 newborns, 2 scientifically confirmed genders 6x4x3x2x1=720

Elsa213:

Whoa e.e

Logic007:

In my opinion,I think 7 different gender sequences

Pixel:

hmm lemme check

Logic007:

1 boy and 5 girls. 2 boys and 4 girls. 3 boys and 3 girls. 4 boys and 2 girls. 5 boys and 1 girls. 6 boys 6 girls

Pixel:

yes you are correct, i did it the wrong way lol

Logic007:

^

rg819616:

which answer choice

Pixel:

7

Pixel:

wait there is n 7

rg819616:

its 720 am i wrong?

Logic007:

or mb \(2^6=64\)

Pixel:

that too^

Logic007:

2 possible sexes for each of 6 babies: \(2*2*2*2*2*2=2^6=64\) Answer would be 64